# LOTFormer: Doubly-Stochastic Linear Attention via Low-Rank Optimal Transport

### v1

@@ Appendix D Implementation Details
Table 7: Hyperparameters for Long Range Arena
Task Learning Rate Sinkhorn Eps Reference Mass Temp
Text 1e41e-4 $5.0$ $0.05$
ListOps 1e31e-3 $0.05$ $8.0$
Retrieval 3e43e-4 $0.1$ $0.1$
Pathfinder 2e42e-4 $0.05$ $0.5$
Image 2e32e-3 $0.05$ $4.0$
D.1 Long Range Arena
We implement our method inside of the Skyformer Chen et al. (2021) codebase and, unless otherwise specified, adopt the same hyperparameters and training schedule as them. For all tasks, we perform learning rate warmup for $5,000$. We use a cosine learning rate scheduler with a minimum learning rate factor of $0.1$. For each task, we sweep learning rate, sinkhorn epsilon, and reference mass temperature. We perform this sweep without a DWC. We select the best hyperparameters based on performance on validation data. For all tasks, we fix number of sinkhorn iterations to 5. We provide the hyperparameters found for each dataset in Table 7. When adding the DWC, we set the kernel size to 3 and utilize a 1d convolution for Text, Retrieval, and ListOps tasks and 2d convolution for Text and Pathfinder. All other hyperparameters are kept fixed.
D.2 ImageNet1K
Table 8: Hyperparameters for ImageNet-1K Classification
Dataset Optimizer Learning Rate (bs=1024) Weight Decay Epochs Warmup
ImageNet-1K AdamW $3\!\times\!10^{-4}$ $5\!\times\!10^{-2}$ $400$ $10$ (Cosine)
@@ D.2 ImageNet1K
We implement our method on top of the official Deit-Tiny training framework and, unless otherwise stated, keep the same training setup. All ImageNet-1K models are trained for 400 epochs using the AdamW optimizer with linear warmup for the first 10 epochs. The base learning rate is $3\mathrm{e}{-4}$ for a global batch size of 1024, and we use weight decay $5\mathrm{e}{-2}$.
D.3 Neural Machine Translation
For our neural machine translation experiments, we adopt the Transformer model from fairseq along with its DiffTransformer counterpart, training both from scratch for 25 epochs. We then fine-tune them alongside other baselines for an additional 10 epochs on the IWSLT’14 dataset [^1] https://github.com/pytorch/fairseq/blob/main/examples/translation/README.md.

### v2

@@ Appendix D Implementation Details
What we change across all experiments.
LOTFormer is evaluated either by training from scratch (ImageNet and LRA) or by conversion of a trained checkpoint (GLUE and translation). In all settings, we only modify bidirectional self attention inside the encoder, and keep token embeddings, FFNs, layer norms, residual wiring, and task specific heads unchanged.
LOTFormer defaults.
Unless a sweep specifies otherwise, we use Sinkhorn iterations T=5T{=}5, learnable reference measure $\mathsf{ref}$ initialized uniform, and the same [CLS] treatment as described in Sec. 4.2. Concretely, [CLS] aggregates via a standard softmax row, while non [CLS] rows use doubly stochastic attention.
Depthwise convolution (DWC).
We treat DWC as an optional token mixing component and report results both with and without it on ImageNet and LRA where applicable. When enabled, we use kernel size $3$ and keep all other hyperparameters fixed.
D.1 ImageNet 1K
Backbones and resolution.
We evaluate at $224^{2}$ resolution on DeiT Tiny, DeiT Small, PVT Tiny, and Swin Tiny. For each backbone, we replace the original bidirectional self attention with LOTFormer while keeping the remaining architecture unchanged.
Training recipe.
@@ D.1 ImageNet 1K
We build on the official DeiT training framework and keep its standard augmentation and regularization recipe unless otherwise stated. All ImageNet models are trained for $400$ epochs using AdamW with linear warmup for the first $10$ epochs and a cosine schedule thereafter. The base learning rate is $3\mathrm{e}{-4}$ for global batch size $1024$, with weight decay $5\mathrm{e}{-2}$.
[CLS] handling and polarization.
We use the dedicated [CLS] softmax aggregator from Sec. 4.2. We also consider [CLS] only polarization as defined in Eq. (6). Table 3 reports controlled ablations for the [CLS] aggregator, [CLS] only polarization, and DWC under matched training settings.
Pivot size and Sinkhorn settings.
For DeiT Tiny, we sweep pivot size $r\in\{4,8,16,32,64\}$ and select the best accuracy cost tradeoff (Table 7). Unless otherwise stated, we use T=5T{=}5 and a learnable reference measure. Additional Sinkhorn hyperparameter sweeps are reported in the ablation section.
Table 8: Training hyperparameters for ImageNet 1K classification.
Dataset Optimizer LR (bs=1024) Weight decay Epochs Warmup
ImageNet 1K AdamW $3\!\times\!10^{-4}$ $5\!\times\!10^{-2}$ $400$ $10$ epochs
D.2 Long Range Arena
Codebase and schedule.
We implement LOTFormer inside the Skyformer (Chen et al., 2021) codebase and follow its standard setup unless otherwise specified. For all tasks, we use 5,0005{,}000 warmup steps and cosine learning rate decay with a minimum factor of $0.1$.
Hyperparameter selection.
For each task, we sweep learning rate, Sinkhorn entropy $\varepsilon$, and reference mass temperature, selecting hyperparameters by validation performance. We fix Sinkhorn iterations to T=5T{=}5 for all tasks. The selected hyperparameters are reported in Table 9.
DWC.
LRA models do not use a [CLS] pooling token. We additionally report LOTFormer with and without the depthwise convolution (DWC) token mixing layer used in PolaFormer. Following PolaFormer’s setup, we use kernel size $3$ with depthwise convolution and keep its associated hyperparameters the same as in PolaFormer. For 1D sequence tasks (Text, Retrieval, ListOps) we use a 1D DWC, and for 2D inputs (Image, Pathfinder) we use a 2D DWC, keeping all other training and LOTFormer hyperparameters fixed.
Table 9: Selected hyperparameters for Long Range Arena.
Task Learning rate Sinkhorn $\varepsilon$ Ref mass temp
Text $1\mathrm{e}{-4}$ $5.0$ $0.05$
ListOps $1\mathrm{e}{-3}$ $0.05$ $8.0$
Retrieval $3\mathrm{e}{-4}$ $0.1$ $0.1$
Pathfinder $2\mathrm{e}{-4}$ $0.05$ $0.5$
Image $2\mathrm{e}{-3}$ $0.05$ $4.0$
D.3 GLUE conversion on fine tuned BERT
Protocol and checkpoints.
We follow the same finetuned conversion protocol as Hedgehog. We start from publicly available fine tuned BERT base uncased checkpoints for each GLUE task, and convert the model by replacing all encoder bidirectional self attention blocks with the candidate attention mechanism.
Attention distillation phase.
For methods that require attention distillation, we train the attention modules for up to 5 epochs with early stopping using AdamW, learning rate $1\mathrm{e}{-2}$, and weight decay 0.
Post conversion training.
After conversion, we train with batch size $8$, AdamW, learning rate $1\mathrm{e}{-5}$, weight decay 0, and a cosine learning rate scheduler for up to 5 epochs. We run this procedure on each GLUE task, using the standard classification objective for all tasks except STS–B, which uses a regression objective.
Reporting.
Table 4 reports per task scores and percent recovery relative to the fine tuned baseline.
D.4 Neural Machine Translation
Backbones and data.
[…]
Two stage evaluation.
First, we pre train the baseline backbone for $25$ epochs. We then perform Plug and Play conversion by swapping encoder bidirectional self attention with LOTFormer, ESPFormer, or Sinkformer, and evaluate with no additional training. Second, we run a fine tune boost phase for $10$ additional epochs. We do not modify decoder causal self attention or encoder decoder cross attention. We report median performance over 4 runs in Table 5.
Training recipe.
We use the standard fairseq IWSLT translation recipe for optimizer, schedule, and regularization. The configuration follows the official fairseq example setup referenced in the main paper.
D.5 Ablations
Sinkhorn and reference measure on ViT.
For the ViT ablations in Table 6, we use ImageNet 100 and a reduced ViT backbone with half hidden dimension. We sweep Sinkhorn iterations $T$, entropy $\varepsilon$, and reference measure choice (fixed versus learnable initialized uniform), and report Top 1 accuracy.
Pivot size on DeiT Tiny.
For the pivot size study in Table 7, we evaluate on ImageNet 1K using DeiT Tiny and sweep $r\in\{4,8,16,32,64\}$. We report Top 1 accuracy and per step forward and backward times measured on a 4 GPU training setup.
Appendix E FlashAttention comparison
Context.
FlashAttention is a fused GPU implementation of exact softmax attention that is optimized for memory traffic via tiling and online softmax, avoiding materializing the full attention matrix and reducing peak memory through recomputation in backward. Our current LOTFormer implementation is written in standard PyTorch operators with autograd, and does not use a custom fused Triton or CUDA kernel.
Disclaimer.
The comparison below is not intended to claim kernel level parity with FlashAttention. It is included as a practical reference point for peak memory under long sequence lengths. Our efficiency claims are algorithmic: for fixed pivot size $r\ll n$, LOTFormer computes attention via low rank OT factors with cost $O(nr)$ and never forms an $n\times n$ attention map. The remaining gap to FlashAttention in peak memory and runtime is dominated by implementation constants such as operator boundaries and autograd saved tensors. Developing a fused implementation that applies FlashAttention style tiling, fusion, and recomputation to the rectangular OT couplings and value mixing is a natural direction for future work and is outside the scope of this work.
Measurement setup.
We measure peak allocated GPU memory for FlashAttention v3 and for our current unfused PyTorch LOTFormer under the same training configuration. We report sequence length $n$, peak memory in GB, and the relative overhead of the current LOTFormer implementation. Further details such as GPU model, dtype, batch size, and precision are provided in the experimental log.
Table 10: Peak GPU memory versus sequence length. FlashAttention v3 is a fused softmax attention kernel. LOTFormer is our current unfused PyTorch implementation. The overhead column reflects current implementation constants, not algorithmic scaling.
Seq len FlashAttention v3 (GB) LOTFormer PyTorch (GB) Overhead
4k 18.2 22.4 +23%
8k 20.8 26.1 +26%
16k 25.9 33.5 +29%
32k 36.2 48.3 +33%
64k 56.8 78.0 +37%
128k 98.0 137.3 +40%
Takeaway.
FlashAttention v3 reflects highly optimized fused kernels for quadratic softmax attention, while our current LOTFormer numbers reflect an unfused PyTorch baseline. The table shows higher constant factor memory for LOTFormer today, even though LOTFormer avoids forming an $n\times n$ attention map and has linear complexity in $n$ for fixed $r$. Closing this constant factor gap is primarily an engineering effort toward fused kernels and recomputation, which we leave to future work.
Appendix F Bidirectional scope and causal masking
Scope.
LOTFormer enforces a doubly stochastic (DS) attention matrix
$A\mathbf{1}=\mathbf{1},\qquad\mathbf{1}^{\top}A=\mathbf{1}^{\top},\qquad A\geq 0,$
which is naturally suited to bidirectional attention (full support).
Causal masking.
Causal self attention restricts the support to a lower triangular set:
$A_{ij}=0\quad\text{for }j>i.$
Under this constraint, the DS constraints force a unique degenerate solution. Indeed, column $n$ can receive mass only from row $n$, so
$\sum_{i=1}^{n}A_{in}=A_{nn}.$
Since DS requires $\sum_{i=1}^{n}A_{in}=1$, we obtain $A_{nn}=1$. Then the row sum constraint $\sum_{j=1}^{n}A_{nj}=1$ together with causality implies $A_{nj}=0$ for all $j<n$. Repeating the same argument for columns $n-1,n-2,\dots,1$ yields $A_{ii}=1$ and $A_{ij}=0$ for $i\neq j$, hence $A=I$. Therefore, DS attention under standard causal masking collapses to the identity and is not compatible with nontrivial autoregressive self attention.
Implication.
LOTFormer targets encoder style bidirectional self attention and cross attention. Extending DS attention to autoregressive settings would require modifying the constraints or the mask structure, which we leave to future work.